# Fast Equivariant Imaging: Acceleration for Unsupervised Learning via Augmented Lagrangian and Auxiliary PnP Denoisers

### v1

@@ 3 Numerical Experiments
We now present a series of experiments in which we verify the proposed FEI scheme. We consider Sparse-view CT Imaging, which was also studied in Chen et al. (2021). We show here that (i) FEI behaves well numerically in both performance and efficiency; (ii) it is better than Equivariant Imaging (EI); (iii) it performs even better when incorporating the PnP Prior; and (iv) PnP-FEI is the fastest unsupervised training algorithm.
Throughout the experiments, we compare our scheme (FEI) with (i) vanilla EI; (ii) PnP-FEI with a pre-trained denoiser, which in practice we consider DnCNN, and BM3D. For these denoising algorithms, we choose the standard deviation of the noise as $\sigma=1.0/\lambda$ ($\lambda$ is the Augemented Lagrangian parameter as presented in Algorithm 3) which are implemented via DeepInv toolbox [^1] https://deepinv.github.io/deepinv/. For a fair comparison with FEI, all the reported results for EI are obtained by directly running the released code of Chen et al. (2021). We use the same U-Net to bulid $G_{\theta}$ as suggested in Chen et al. (2021) for EI, FEI and PnP-FEI. Specifically, $G_{\theta}$ is a four-depth residual U-Net (Ronneberger et al., 2015) with a symmetric encoder-decoder and long skip connections. Each encoder comprises two $3\times 3$ Convolution-BatchNorm-ReLU blocks with identity shortcuts, followed by $2\times 2$ max pooling to double channels $[64,128,256,512]$. The decoder mirrors the encoder in a refinement manner, and a final $1\times 1$ convolution restores the output. For the measurement consistency loss, we choose $\ell_{2}$ in this experiment for simplicity.
For the sparse-view CT task, when updating $\theta$, we use the Adam optimizer with an initial learning rate of 0.001 in all methods. The weight decay is $1\times 10^{-8}$. We trained the networks for 10,000 epochs, keeping the learning rate constant for the first 5000 epochs, and then shrinking it by a factor of $0.5\times(1+\cos(\frac{\#\text{epoch}}{10000}\cdot\pi))$. We verify our model with equivariance strength $\alpha=100$ as suggested in Chen et al. (2021). Other parameters of the FEI scheme are set as follows: Lagrangian multiplier $\lambda=1$; NAG momentum factor $\beta=0.1$; NAG step size $\eta=0.01$; NAG iterations $J=10$. The complete NAG algorithm is presented in Algorithm 2. The gradient descent step size in PnP-FEI is $\gamma=0.01$. We also use a moving average to smooth the reconstructed images for better reconstruction and smooth PSNR curves, the same as Sun et al. (2021).
@@
In order to better understand the boosts of proposed FEI scheme for sparse-view CT Imaging, we show the PSNR curves with respect to both iterations and times in Figure 1, and Table 3 shows the PSNR in the last iteration in Figure 1. In Figure 1 we can observe that the PnP-FEI is able to provide order-of-magnitude (10$\geq$ 10$\times$ 10 ×) acceleration over standard EI training.
@@
[…]
tablePSNR on training dataset at Figure 1 for Sparse-view CT reconstruction.
@@
FBP 27.33 $\pm$ 0.64
EI 32.88 $\pm$ 1.21
FEI 33.20 $\pm$ 1.42
PnP-FEI (BM3D) 33.10 $\pm$ 1.22
PnP-FEI (DnCNN) 33.56 $\pm$ 1.21
EQPnP-FEI (DnCNN) 33.44 $\pm$ 1.18
Supervised training 33.71 $\pm$ 1.37
Table 1: (Generalization performance) Averaged PSNR of 10 test images (index 11 to 20) for Sparse-view CT reconstruction.
@@
Figure 3: Test on another unseen CT100 sample with Radon scan number is 40 (a test on out-of-distribution generalization).
@@ 4 Conclusion
We propose FEI and its variant PnP-FEI for accelerating unsupervised training of deep neural networks for imaging inverse problems without the need of ground-truth data. We observe that our framework is capable of significantly accelerating the state-of-the-art unsupervised training scheme Equivariant Imaging proposed by (Chen et al., 2021), along with improved generalization performance. We believe that this work is crucial for the area of deep learning for computational imaging, via making the unsupervised training truly practical. Our algorithmic framework (Augmented Lagrangian plus PnP) can be easily extended to more recent variants of EI such as Robust Equivariant Imaging (REI) (Chen et al., 2022), Multi-Operator Imaging (MOI) (Tachella et al., 2022), and our Sketched Equivariant Imaging (SkEI) (Xu et al., 2024), etc.

### v2

@@ 3 Numerical Experiments
This section describes the quantitative and qualitative experiments conducted to evaluate the performance of the proposed method.
Framework Optimization Backbone           Description
EI Adam Vanilla EI framework, which use back-propagation and
end-to-end training strategy.
FEI-Option1 Adaptive HQS HQS with accelerated gradient descent to approximately
minimize Latent-reconstruction step subproblem,
while Adam for online Pseudo-supervision, Algorithm 1.
FEI-Option2 Adaptive Linearized ADMM The difference from FEI-Option1 is the change of splitting
style to linearized ADMM, with gradient step for Latent
-reconstruction, while Adam for online Pseudo-supervision,
Algorithm 3.
PnP-FEI Adaptive Linearized ADMM The difference from FEI-Option2 is the use of pre-trained
denoiser for Latent-reconstruction step instead of plain
gradient step, Algorithm 4.
EQPnP-FEI Adaptive Linearized ADMM The difference from PnP-FEI is that we use equivariant
plug-and-play to perform the Latent-reconstruction step,
Algorithm 5.
Table 1: Unsupervised learning algorithms in our experiments.
Throughout the experiments, we compare our scheme (FEI) with (i) vanilla EI; (ii) PnP-FEI with a pre-trained denoiser, which in practice we consider DnCNN and BM3D. For these denoising algorithms, we choose the standard deviation of the noise as $\sigma=1.0/\lambda$ ($\lambda$ is the augmented Lagrangian parameter as presented in Algorithm 4) which are implemented via DeepInv toolbox [^1] https://deepinv.github.io/deepinv/. For a fair comparison with FEI, all the reported results for EI are obtained by directly running the released code of Chen et al. (2021). We use the same U-Net to bulid $G_{\theta}$ as suggested in Chen et al. (2021) for EI, FEI and PnP-FEI. Specifically, $G_{\theta}$ is a four-depth residual U-Net (Ronneberger et al., 2015) with a symmetric encoder-decoder and long skip connections. Each encoder comprises two $3\times 3$ Convolution-BatchNorm-ReLU blocks with identity shortcuts, followed by $2\times 2$ max pooling to double channels $[64,128,256,512]$. The decoder mirrors the encoder in a refinement manner, and a final $1\times 1$ convolution restores the output. For the measurement consistency loss, for simplicity, we choose $\ell_{2}$ in this experiment.
3.1 Sparse-view CT
For the sparse-view CT task, when updating $\theta$, we use Adam optimizer with an initial learning rate of 0.001 in all methods. The weight decay is $1\times 10^{-8}$. We trained the networks for 10,000 epochs, keeping the learning rate constant for the first 5000 epochs, and then shrinking it by a factor of $0.5\times(1+\cos(\frac{\#\text{epoch}}{10000}\cdot\pi))$. We verify our model with equivariance strength $\alpha=100$ as suggested in Chen et al. (2021). Other parameters of the FEI scheme are set as follows: Lagrangian multiplier $\lambda=1$; NAG momentum factor $\beta=0.1$; NAG step size $\eta=0.01$; NAG iterations $J=10$. The complete NAG algorithm is presented in Algorithm 2. The gradient descent step size in FEI-Option2 and PnP-FEI is $\gamma=0.01$. We also use a moving average to smooth the reconstructed images for better reconstruction and smooth PSNR curves, the same as Sun et al. (2021).
@@ 3.1 Sparse-view CT
In order to better understand the boosts of proposed FEI scheme for sparse-view CT Imaging, we show the PSNR curves with respect to both iterations and times in Figure 1, and Table 2 shows the PSNR in the last iteration in Figure 1. In Figure 1 we can observe that the PnP-FEI is able to provide order-of-magnitude ($\geq 10\times$) acceleration over standard EI training.
@@
Table 2: Averaged PSNR on training dataset at Figure 1 for Sparse-view CT reconstruction, * is baseline model.
@@
FBP 27.98 $\pm$ 1.23
EI 33.23 $\pm$ 1.55
FEI-Option1 33.93 $\pm$ 1.97
FEI-Option2 33.92 $\pm$ 1.77
PnP-FEI (BM3D) 33.95 $\pm$ 1.93
PnP-FEI (DnCNN) 34.06 $\pm$ 1.21
EQPnP-FEI (DnCNN) 34.12 $\pm$ 1.73
Supervised training 34.61 $\pm$ 2.12
Table 3: (Generalization performance) Averaged PSNR of 10 test images (index 11 to 20) for Sparse-view CT reconstruction, * is baseline model.
@@
Figure 3: Test on unseen CT100 sample with Radon scan number is 40 (a test on out-of-distribution generalization).
3.2 Inpainting
For the inpainting task, when updating $\theta$, we use the Adam optimizer with an initial learning rate of 0.001 in all methods. The weight decay is $1\times 10^{-8}$. We trained the networks for 5,000 epochs, keeping the learning rate constant for the first 4,000 epochs, and then shrinking it by a factor of $0.5\times(1+\cos(\frac{\#\text{epoch}}{5000}\cdot\pi))$. We verify our model with equivariance strength $\alpha=1$ as suggested in Chen et al. (2021). Other parameters of the FEI scheme are set as follows: Lagrangian multiplier $\lambda=0.4$; NAG momentum factor $\beta=0.9$; NAG step size $\eta=0.09$; NAG iterations $J=10$. The complete NAG algorithm is presented in Algorithm 2. The size of the gradient descent step in FEI-Option2 and PnP-FEI is $\gamma=1.0$. We also use a moving average to smooth the reconstructed images for better reconstruction and smooth PSNR curves, the same as Sun et al. (2021).
The mask rate for the inpainting task is $0.6$, which means 60% of pixels are randomly removed. Here we choose random shift as the transformation. The dataset we used here is Urban100 (Huang et al., 2015) which is the same as Chen et al. (2021). For each image, we cropped a $512\times 512$ pixel area in the center and then resized it to $256\times 256$ for the ground truth image. The first 90 measurements are for training, while the last 10 images are for testing. We again observe significant benefit of FEI over standard EI training.
Figure 4: PSNR and MSE curves with respect to iteration during the training of inpainting reconstruction. We observe a significant acceleration can be achieved by FEI over EI (we fine-tuned EI with optimal algorithmic parameters for fastest convergence, but we did not fine-tune our own methods, so there is likely even more room for acceleration).
Methods PSNR (dB)
*EI 24.18 $\pm$ 2.59
FEI-Option1 26.50 $\pm$ 2.74
FEI-Option2 25.11 $\pm$ 2.57
PnP-FEI (DnCNN) 25.46 $\pm$ 2.83
EQPnP-FEI (DnCNN) 25.44 $\pm$ 2.83
Table 4: PSNR on training dataset at Figure 4 for inpainting reconstruction, * is baseline model.
Methods PSNR (dB)
FBP 2.80 $\pm$ 1.66
EI 21.19 $\pm$ 1.95
FEI-Option1 22.00 $\pm$ 2.10
FEI-Option2 21.51 $\pm$ 2.04
PnP-FEI (DnCNN) 21.39 $\pm$ 1.90
EQPnP-FEI (DnCNN) 21.81 $\pm$ 1.97
Supervised training 22.32 $\pm$ 2.30
Table 5: (Generalization performance) Averaged PSNR of 10 test images for inpainting reconstruction, * is baseline model.
Figure 5: Test on another unseen Urban100 sample with mask rate is 0.6.
@@ 4 Conclusion
We propose FEI and its variant PnP-FEI for accelerating unsupervised training of deep neural networks to imaging inverse problems without the need of ground-truth data. We observe that our framework is capable of significantly accelerating the state-of-the-art unsupervised training scheme Equivariant Imaging proposed by (Chen et al., 2021), along with improved generalization performance. We believe that this work is crucial for the area of deep learning for computational imaging, via making the unsupervised training truly practical. Our algorithmic framework (Augmented Lagrangian plus PnP) can be easily extended to more recent variants of EI such as Robust Equivariant Imaging (REI) (Chen et al., 2022), Multi-Operator Imaging (MOI) (Tachella et al., 2022), and our Sketched Equivariant Imaging (SkEI) (Xu et al., 2024), etc.